# Critical Iterative Denoising A Discrete Generative Model Applied to Graphs

### v1

@@ 4.3 Implementation and Sampling
Since $p_{\alpha_{t}}(a)$ does not depends on $\hat{Z}_{1|t}$, $\mathbb{E}_{\hat{Z}_{1|t}}p(a|\hat{Z}_{1|t})=p_{\alpha_{t}}(a)=\alpha_{t}$, we actually cast the Critic as a predictor of the residual logit with respect to the true $\alpha_{t}$:
$p_{\phi}(a|\hat{Z}_{1|t})=\sigma(f_{\phi}(\hat{Z}_{1|t},\alpha_{t})+\sigma^{-1}(\alpha_{t})),$ (18)
where $\sigma$ is the sigmoid function and $f_{\phi}(\hat{Z}_{1|t},\alpha_{t})$ is a GNN. The Critic operates over the denoiser’s outputs; we train it post hoc with a fixed denoiser. We can leverage any denoiser implementing mask diffusion and no retraining is needed.
During inference, we are interested in $\hat{a}_{t+\Delta_{t}}$ rather than $\hat{a}_{t}$. We therefore use the approximation $\hat{\alpha}_{t+\Delta_{t}}\approx\sigma(f_{\phi}(\hat{Z}_{1|t},\alpha_{t})+\sigma^{-1}(\alpha_{t+\Delta_{t}}))$.
In summary a denoising step involves: (1) sampling $\hat{Z}_{1|t}$ from the Denoiser, getting $\hat{\alpha}_{t+\Delta_{t}}$ from the Critic, masking back using Equation 6, replacing $\alpha_{t}$ with $\hat{\alpha}_{t+\Delta_{t}}$.
@@ 5 Evaluation
We evaluate our model on molecular and synthetic graph datasets.
For molecular data, we use the Qm9 and Zinc250k datasets. The Qm9 dataset contains 133,885 molecules with up to 9 atoms of 4 types, while the Zinc250k dataset consists of 250,000 molecular graphs with up to 38 heavy atoms of 9 types. For generic graphs, we run experiments on the Planar and the Stochastic Block Model (SBM) datasets. Both contain 200 unattributed graphs, with 64 nodes and up to 200 nodes, respectively. Visualizations of generated molecules and graphs are available in Appendix F.
Our objective is to evaluate our proposed method in direct comparison with Discrete Diffusion Models (DDM). To ensure a rigorous comparison, we train two denoisers, one using the marginal distribution and the other using the mask distribution. We use these denoisers to compare the two sampling producures: Discrete Diffusion and Iterative Densoising. Additionally, we train a Critic associated with the model using the mask distribution. Specifically, the Marginal DDM and Marginal ID share the same denoiser, while the Mask DDM, Mask ID, and Mask CID also share the same denoiser (architecture and parameters).
This setup ensures a controlled and fair ablation study, isolating the impact of our iterative denoising approach while maintaining identical model architectures and training conditions across all comparisons. We provide the technical experimental details and implementation in Appendix D.
We also provide baselines results from various diffusion models; GDSS (Jo et al., 2022) is a continuous diffusion model, DruM (Jo et al., 2024) is a diffusion bridge model, and DiGress (Vignac et al., 2023) is a Discrete Diffusion model. DiGress uses the marginal distribution as noise. It differs from our Marginal DDM only by the network architecture, and hyperparameters. Results are taken from Jo et al. (2024). Importantly, these baseline results were obtained using 1000 function evaluation steps during generation, whereas we use 500 steps in our experiments.
@@ 5.1 Molecule Generation
Table 1: Results on Qm9 and Zinc250k datasets.
@@
Marg. ID $99.67$ $\mathbf{1.04}$ $\mathbf{0.50}$ $99.50$ $\mathbf{2.06}$ $\mathbf{2.01}$
Mask ID $96.43$ $1.40$ $1.80$ $93.85$ $11.08$ $9.05$
@@
The results are presented in Table 1. We highlight the best results among our experiments. We also highlight the best baseline if better than our models. We generate 10,000 samples and compare them with 10,000 test samples. Results are averaged over 5 sampling runs. Standard deviations and additional metrics (e.g., uniqueness and novelty) are included in Appendix E.
We observe that our Iterartive Denoising (ID) models systematically outperform their Discrete Diffusion (DDM) counterparts using the same denoiser on both datasets and this with both noise distributions, i.e. marginal and mask. Furthermore, the results confirm our analysis of compounding denoising error, particularly affecting mask-based discrete diffusion, which performs poorly across both datasets. ID and CID address the compounding error successfully. Our experiments confirm that the marginal distribution is a more effective noise distribution in graph modeling compared to the mask distribution. Even if the compounding denoising error is less pronounced with the marginal distribution, we still observe that our Iterative Denoising method outperforms consistently the marginal prior discrete diffusion baselines (DDM, Marg. DDM) by a large margin. Finally, we observe that the Critic further improves the generative performance. Notably, our Critical Iterative Denoising method reaches a high-level validity rate, with only a few invalid graphs, specifically on Zinc250k invalid molecules are 50 less frequent in our model than in the best baseline model (DruM).
@@ 5.2 Generic graphs
For generic graphs, we follow the evaluation procedure established by Martinkus et al. (2022) and used by Jo et al. (2022) and Jo et al. (2024), using an 80/20 train-test split, with 20% of the training data reserved for validation.
Table 2: Results on Planar and SBM datasets
@@
Marg. ID $7.62$ $\mathbf{91.3}$ $\mathbf{5.93}$ $\mathbf{63.5}$
Mask ID $8.72$ $67.0$ $15.05$ $17.5$
@@
We evaluate graph similarity using the Spectral Maximum Mean Discrepancy (Spect.), which compares the spectra of the graph Laplacian between generated and real graphs. Unlike other metrics that rely on local structural properties, such as node degrees or clustering coefficients, the Spectral MMD captures graph structures at any level. Additionally, we report the rate of Valid, Unique, and Novel graphs (V.U.N.). For the Planar dataset, a valid graph must be both planar and connected. For the Stochastic Block Model dataset, validity indicates that the generated graph is likely to follow the block model distribution used to generate the training data (i.e., an intra-community edge density of 0.3 and an inter-community edge density of 0.005). Uniqueness represent the fraction of unique graphs among the generated samples, and novelty the fraction of unique graphs that do not appear in the training set. In practice, all generated graphs in our experiments were both unique and novel, meaning that the V.U.N. metric actually reports the validity rate. We present our results in Table 2. We report the average over 5 sampling comparing the test set and a generative sample batch of same size. Standard deviation and complementary results including MMDs based on degree, clustering coefficient, and orbits counts on Appendix E.
The analysis of general graphs aligns with our findings on molecular graphs, further reinforcing our conclusions. Our Iterative Denoising models consistently outperform their Discrete Diffusion counterparts, demonstrating their effectiveness across different graph types. Results also confirm that mask-based models are not well suited for graph generation. In particular, on the SBM dataset, their results are barely significant due to their performances .
While the Critic improves the spectral MMD on the Planar dataset, we observe that its contribution is less effective for general graphs compared to molecular graphs. We hypothesize that this is due to deviations from the data distribution being less localized in generic graphs, making it more difficult to discriminate between graph elements. With the marginal distribution, our Iterative Denoising model presents results that are significantly better than the Discrete Diffusion baseline.
@@ 5.3 Ablation
We assess the impact of the Number of Function Evaluations (NFE)—i.e., the number of denoising steps during sampling—on model performance. Specifically, we investigate how the Critic influences the required NFE.
We generate graphs with varying NFEs in $\{16,32,64,128,256,512\}$ and evaluate performance on the Zinc250k dataset for Marginal DDM, Marginal ID, and Mask CID. The validity rate is reported in Figure 1, while additional figures for other metrics and full numerical results are provided in Appendix E due to space limitation.
Our findings indicate that the Critic significantly reduces the required NFEs to achieve a high validity rate, reaching over 99% validity in just 32 steps. Furthermore, in the lowest-NFE regime (16 steps), Mask CID consistently outperforms all other models across all metrics, demonstrating its effectiveness in reducing the number of denoising steps.
@@ 6 Conclusion
Discrete diffusion models suffer from the Compounding Denoising Error issue, which leads to error accumulation during sampling and negatively affects their performance. In this paper, we address this issue by removing the time dependency in the noising process. The resulting method does not rely on a diffusion process. We further improve upon our method, and introduce a Critic, which steers the distribution of generated elements towards the data distribution, improving sample quality.
Our experimental results demonstrate the effectiveness of our method. Our Iterative Denoising model consistently outperforms its corresponding denoising model. Moreover, our Critic-based model systematically improves generative performance over the corresponding Mask Iterative Denoising model without the Critic.
@@ A.3 Denoising
$\begin{split}p(z_{s}|z_{t})&=\sum_{z}p_{s|1}(z|z_{t},z_{1})p_{1|t}^{\theta}(z|Z_{t})\\ &=\sum_{z}p_{s|1}(z|x_{1})p_{1|t}^{\theta}(z|Z_{t})\\ &=\sum_{z}p_{s|1}(z|x_{1})p_{1|t}^{\theta}(z|Z_{t})\\ &=\sum_{z}\sum_{a\in\{0,1\}}p_{s|1}(z|x_{1},a_{s})p_{s}(a)p_{1|t}^{\theta}(z|Z_{t})\\ &=\sum_{z}(q_{1}(z)\alpha_{s}+q_{0}(z)(1-\alpha_{s}))p_{1|t}^{\theta}(z|Z_{t})\\ &=\sum_{z}(p_{1|t}^{\theta}(z|Z_{t})q_{1}(z)\alpha_{s})+\sum_{z}p_{1|t}^{\theta}(z|Z_{t})q_{0}(z)(1-\alpha_{s}))\\ &=(p_{1|t}^{\theta}(z|Z_{t})\alpha_{s})+q_{0}(z)(1-\alpha_{s}))\sum_{z}p_{1|t}^{\theta}(z|Z_{t})\\ &=\alpha_{s}p_{1|t}(z|z_{t})+(1-\alpha_{s})q_{0}(z)\end{split}$ (24)
@@
The proof is strongly inspired by Goodfellow et al. (2014).
$\begin{split}{\mathcal{L}}_{\phi}&=-\mathbb{E}_{t,\hat{Z}_{1|t}}p_{\phi}(a|\hat{Z}_{1|t})\\ &=-\alpha_{t}\mathbb{E}_{\hat{Z}_{1|t}\sim p_{data}}p_{\phi}(a|\hat{Z}_{1|t})\\ &\quad-(1-\alpha_{t})\mathbb{E}_{\hat{Z}_{1|t}\sim p_{pred}}p_{\phi}(a|\hat{Z}_{1|t})\end{split}$ (26)
From Equation 13, we have:
$\begin{split}\mathbb{E}_{\hat{Z}_{1|t}}p_{\phi}(a|\hat{Z}_{1|t})&=\alpha_{t}\mathbb{E}_{\hat{Z}_{1|t}\sim p_{data}}p_{\phi}(a|\hat{Z}_{1|t})+(1-\alpha_{t})\mathbb{E}_{\hat{Z}_{1|t}\sim p_{pred}}p_{\phi}(a|\hat{Z}_{1|t})\\ &=\sum_{z\in{\mathcal{Z}}}\alpha_{t}p_{data}(\hat{z}_{1|t})p_{\phi}(a|\hat{Z}_{1|t})+(1-\alpha_{t})p_{pred}(\hat{z}_{1|t})p_{\phi}(a|\hat{Z}_{1|t})\end{split}$ (27)
@@
Hence, Equation 27 reaches its maximum at:
@@
Appendix B Generative Graph Modeling: Related Works
One of the main challenges in graph modeling arises from the multiple possible representations of a single graph, due to the $n!$ possible node permutations. This complexity has led to two main approaches: sequential models and equivariant models.
Sequential models generate graphs auto-regressively, progressively adding nodes, edges, or subgraphs (You et al., 2018; Shi et al., 2020; Luo et al., 2021; Liao et al., 2019; Kong et al., 2023; Gómez-Bombarelli et al., 2018; Kusner et al., 2017; Goyal et al., 2020; Jin et al., 2018, 2020). Conversely, Equivariant models tackle the node permutation problem by ensuring a unique computational graph regardless of the specific instantiation of the graph structure. These models have been developed under various generative frameworks, including GANs (Krawczuk et al., 2021; Martinkus et al., 2022), normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020; Liu et al., 2019), and vector-quantized auto-encoders (Boget et al., 2024; Nguyen et al., 2024).
Recently, equivariant denoising models have employed continuous denoising techniques, such as score-based models (Yang et al., 2019; Jo et al., 2022), diffusion bridges (Jo et al., 2024), and flow matching (Eijkelboom et al., 2024). However, these continuous relaxations disrupt the discrete structure of graphs during the noising process. Discrete diffusion models in discrete time (Haefeli et al., 2022; Vignac et al., 2023) and continuous time (Xu et al., 2024) have achieved significant improvements in generating small graphs. Additionally, Qin et al. (2024a) recently adapted the discrete flow matching framework for graph modeling.

### v2

@@ 4.3 Implementation and Sampling
Since $p_{\alpha_{t}}(a)$ (Equation 11) does not depend on $\hat{Z}_{1|t}$, $\mathbb{E}_{\hat{Z}_{1|t}}p(a\mid\hat{Z}_{1|t})=p_{\alpha_{t}}(a)=\alpha_{t}$, we actually cast the Critic as a predictor of the residual logit with respect to the true $\alpha_{t}$:
$p_{\phi}(a\mid\hat{Z}_{1|t})=\sigma(f_{\phi}(\hat{Z}_{1|t},\alpha_{t})+\sigma^{-1}(\alpha_{t})),$ (18)
where $\sigma$ is the sigmoid function and $f_{\phi}(\hat{Z}_{1|t},\alpha_{t})$ is a GNN. The Critic operates on the outputs of a fixed denoiser and is trained post hoc, requiring no retraining of the denoiser itself. We can leverage any denoiser implementing mask diffusion and no retraining is needed.
During inference, we are interested in $\hat{\alpha}_{t+\Delta_{t}}$ rather than $\hat{\alpha}_{t}$. We therefore use the approximation $\hat{\alpha}_{t+\Delta_{t}}\approx\sigma(f_{\phi}(\hat{Z}_{1|t},\alpha_{t})+\sigma^{-1}(\alpha_{t+\Delta_{t}}))$.
In summary, each denoising step involves: 1. sampling $\hat{Z}_{1|t}$ from the Denoiser, 2. computing $\hat{\alpha}_{t+\Delta_{t}}$ via the Critic, 3. re-noising using Equation 6, with $\alpha_{t}$ replaced by $\hat{\alpha}_{t+\Delta_{t}}$.
@@ 5 Evaluation
We evaluate our model on both molecular and synthetic graph datasets.
For molecular data, we use the QM9 and ZINC250k datasets. The QM9 dataset contains 133,885 molecules with up to 9 atoms of 4 types, while the ZINC250k dataset consists of 250,000 molecular graphs with up to 38 heavy atoms of 9 types. For generic graphs, we run experiments on the Planar and Stochastic Block Model (SBM) datasets. Both contain 200 unattributed graphs with 64 nodes and up to 200 nodes, respectively. Visualizations of generated molecules and graphs are available in Appendix G.
Our objective is to evaluate the proposed method through direct comparison with Discrete Diffusion Models (DDMs). To ensure a rigorous comparison, we train two denoisers, one using the marginal distribution and the other using the mask distribution. We use these denoisers to compare the two sampling procedures: Discrete Diffusion and Simple Iterative Denoising. Additionally, we train a Critic for the model using the mask distribution. Specifically, the Marginal DDM and Marginal SID share the same denoiser, while the Mask DDM, Mask SID, and Mask CID use a another shared denoiser with identical architecture and parameters.
This setup ensures a controlled and fair ablation study, isolating the impact of our iterative denoising approach while maintaining identical model architectures and training conditions across all comparisons. We provide the technical experimental details and implementation in Appendix E.
We report baseline results from various diffusion models; GDSS (Jo et al., 2022) is a continuous diffusion model, DruM (Jo et al., 2024) is a diffusion bridge model, and DiGress (Vignac et al., 2023) is a Discrete Diffusion model. DiGress uses the marginal distribution as noise. It differs from our Marginal DDM only by the network architecture, and hyperparameters. Baseline results are reported from Jo et al. (2024). Importantly, these baseline results were obtained using 1000 function evaluation steps during generation, whereas we use only 500 steps in our experiments. We report additional computational cost details for our model and the baselines in Appendix E.3.
@@ 5.1 Molecule Generation
Table 1: Molecule generation results on QM9 and ZINC250k.
@@
Marg. SID $99.67$ $\mathbf{1.04}$ $\mathbf{0.50}$ $99.50$ $\mathbf{2.06}$ $\mathbf{2.01}$
Mask SID $96.43$ $1.40$ $1.80$ $93.85$ $11.08$ $9.05$
@@
Table 1 presents the results. The best-performing models among our experiments are bolded. We also bold the best baseline if it outperforms ours. We generate 10,000 samples, compared against 10,000 test molecules. Results are averaged over five independent sampling runs. Standard deviations and additional metrics, including uniqueness and novelty, are reported in Appendix F.
We observe that our Simple Iterative Denoising (SID) models consistently outperform their Discrete Diffusion Model (DDM) counterparts using identical denoisers on both datasets and this across both noise distributions: marginal and mask. Furthermore, the results confirm our analysis of compounding denoising error, particularly affecting mask-based discrete diffusion, which performs poorly across both datasets. SID and CID address the compounding error successfully. Our experiments confirm that the marginal distribution is a more effective noise distribution in graph modeling compared to the mask distribution. While compounding denoising errors are less severe with the marginal distribution, SID still consistently outperforms marginal DDM baselines by a wide margin. Finally, we observe that the Critic further improves the generative performance. Notably, on ZINC250k, our Critical Iterative Denoising model reaches a high-level validity rate, producing 50 times less invalid molecules than the best-performing baseline, DruM.
@@ 5.2 Generic Graphs
For generic graphs, we follow the evaluation protocol introduced by Martinkus et al. (2022) and adopted by Jo et al. (2022) and Jo et al. (2024), using an 80/20 train–test split with 20% of the training data reserved for validation.
Table 2: Graph generation results on Planar and SBM datasets.
@@
Marg. SID $7.62$ $\mathbf{91.3}$ $\mathbf{5.93}$ $\mathbf{63.5}$
Mask SID $8.72$ $67.0$ $15.05$ $17.5$
@@
We evaluate graph similarity using the Spectral Maximum Mean Discrepancy (Spect.), which compares the spectra of the graph Laplacian between generated and real graphs. Unlike other metrics that rely on local structural properties, such as node degrees or clustering coefficients, the Spectral MMD captures graph structures at any level. Additionally, we report the proportion of Valid, Unique, and Novel graphs (V.U.N.). For the Planar dataset, a valid graph must be both planar and connected. For the Stochastic Block Model dataset, validity indicates that the generated graph is likely to follow the block model distribution used to generate the training data (i.e., an intra-community edge density of 0.3 and an inter-community edge density of 0.005). Uniqueness represents the fraction of unique graphs among the generated samples, and novelty the fraction of unique graphs that do not appear in the training set. In practice, all generated graphs in our experiments were both unique and novel; thus, V.U.N. effectively measures validity. Table 2 reports the results. We average over five sampling runs, comparing the test set to a generated sample batch of equal size. Standard deviation and complementary metrics including MMDs based on degree, clustering coefficient, and orbits counts are reported in Appendix F.
The results on general graphs align with our findings on molecular graphs, further reinforcing our conclusions. Our Simple Iterative Denoising models consistently outperform their Discrete Diffusion counterparts, demonstrating their effectiveness across different graph types. Results also confirm that mask-based models are not well suited for graph generation. In particular, on the SBM dataset, their results are barely significant due to their performances.
While the Critic improves spectral MMD on the Planar dataset, its contribution is less pronounced for generic graphs than for molecular graphs. We hypothesize that this is due to deviations from the data distribution being less localized in generic graphs, making it more difficult to discriminate between graph elements. Using the marginal distribution, our Simple Iterative Denoising model achieves results significantly superior to the Discrete Diffusion baseline.
@@ 5.3 Ablation
We assess the impact of the Number of Function Evaluations (NFE)—i.e., the number of denoising steps during sampling—on model performance. Specifically, we investigate how the Critic affects the number of steps required to achieve high-quality generation.
We generate graphs with varying NFEs in $\{16,32,64,128,256,512\}$ and evaluate performance on the ZINC250k dataset for Marginal DDM, Marginal SID, and Mask CID. Validity is reported in Figure 1, while additional plots and full numerical results are provided in Appendix F.
Our findings indicate that the Critic substantially reduces the number of denoising steps needed to achieve high validity, Mask CID reaching over 99% validity in just 32 steps. Notably, in the low-NFE regime (16 steps), Mask CID consistently outperforms all other models across all metrics, demonstrating its effectiveness in accelerating inference without sacrificing quality.
@@ 6 Conclusion
Discrete diffusion models suffer from the Compounding Denoising Error issue, which leads to error accumulation during sampling and negatively affects their performance. We address this issue by introducing an assumption of conditional independence between intermediate noisy states. Thus, the resulting method does not rely on a diffusion process. We further improve upon our method, and introduce a Critic, which steers the distribution of generated elements towards the data distribution, improving sample quality.
Our experimental results demonstrate the effectiveness of our method. Our Simple Iterative Denoising models consistently outperform their corresponding discrete diffusion models. Moreover, our mask model with Critic systematically improves generative performance over the corresponding Simple Iterative Denoising mask model without the Critic.
@@
Acknowledgments
I am deeply grateful to Prof. Alexandros Kalousis, whose support has been invaluable—far beyond what co-authorship could acknowledge.
I acknowledge the financial support of the Swiss National Science Foundation within the LegoMol project (grant no. 207428). The computations were performed at the University of Geneva on ”Baobab”, ”Yggdrasil”, and ”Bamboo” HPC clusters.
@@ A.3 Denoising
Given the random variable $a_{t}\sim p_{\alpha_{t}}(a)=\text{Bernoulli}(a;\alpha_{t})$ (see Section 4.1), we have: $p_{s|1}(z|z_{1},a_{s})=q_{0}(z)\>\text{if}\>z=0$, and $p_{s|1}(z|z_{1},a_{s})=\delta_{z_{1}}(z)$, otherwise.
$\begin{split}p_{s|t}(z|z_{t})&=\sum_{z_{1}}p_{s|1}(z|z_{t},z_{1})p_{1|t}(z_{1}|Z_{t})\\ &=\sum_{z_{1}}p_{s|1}(z|z_{1})p_{1|t}(z_{1}|Z_{t})\\ &=\sum_{z_{1}}\sum_{a}p_{\alpha_{s}}(a)p_{s|1}(z|z_{1},a_{s})p_{1|t}(z_{1}|Z_{t})\\ &=\sum_{z_{1}}\left(\alpha_{s}\delta_{z_{1}}(z)+q_{0}(z)(1-\alpha_{s})\right)p_{1|t}(z_{1}|Z_{t})\\ &=\sum_{z_{1}}p_{1|t}(z_{1}|Z_{t})\alpha_{s}\delta_{z_{1}}(z)+\sum_{z_{1}}p_{1|t}(z_{1}|Z_{t})(1-\alpha_{s})q_{0}(z)\\ &=\alpha_{s}p_{1|t}(z|Z_{t})+(1-\alpha_{s})q_{0}(z)\end{split}$ (24)
@@
The proof is inspired by Goodfellow et al. (2014).
For a single element, from Equation 13, we have:
$\begin{split}{\mathcal{L}}_{\phi}&=-\mathbb{E}_{t,\hat{Z}_{1|t}}\log\left(p_{\phi}(a|\hat{Z}_{1|t})\right)\\ &=-\alpha_{t}\mathbb{E}_{\hat{Z}_{1|t}\sim p_{data}}\log\left(p_{\phi}(a|\hat{Z}_{1|t})\right)-(1-\alpha_{t})\mathbb{E}_{\hat{Z}_{1|t}\sim p_{pred}}\log\left(1-p_{\phi}(a|\hat{Z}_{1|t})\right)\\ &=-\sum_{z_{1}}\alpha_{t}p_{data}(\hat{z}_{1|t})\log\left(p_{\phi}(a|\hat{Z}_{1|t})\right)-\sum_{z_{1}}(1-\alpha_{t})p_{pred}(\hat{z}_{1|t})\log\left(1-p_{\phi}(a|\hat{Z}_{1|t})\right)\\ &=-\sum_{z_{1}}\left(\alpha_{t}p_{data}(\hat{z}_{1|t})\log\left(p_{\phi}(a|\hat{Z}_{1|t})\right)+(1-\alpha_{t})p_{pred}(\hat{z}_{1|t})\log\left(1-p_{\phi}(a|\hat{Z}_{1|t})\right)\right)\end{split}$ (26)
@@
Hence, Equation 26 reaches its minimum at:
@@
Appendix B Comparing Simple Iterative Denoising with Discrete Diffusion and Discrete Flow Matching
In Table 3, we compare Simple Iterative Denoising with Discrete Diffusion and Discrete Flow Matching. For Discrete Diffusion, we follow the framework of Austin et al. (2021), which uses uniform and mask (absorbing state) noise , and its extension to the marginal setting by Vignac et al. (2023). For Discrete Flow Matching, we adopt the standard formulation from Gat et al. (2024), which uses independent coupling and convex interpolants. We use the following shorthand notations: $s=t+\Delta_{t}$ and $D_{t}=\Delta_{t}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}$.
Table 3: Comparing Simple Iterative Denoising, Discrete Diffusion, and Discrete Flow Matching.
Simple Iterative Denoising Discrete Diffusion Discrete Flow Matching
Noising Conditional indep. given $z_{1}$ Markovian Markovian
$p_{t|s,1}(z|z_{s},z_{1})$ $p_{t|1}(z|z_{1})$ $p_{t|s}(z|z_{s})$ $p_{t|s}(z|z_{s})$
$q_{t|1}(z|z_{1})$ $\alpha_{t}\delta_{z_{1}}(z)+(1-\alpha_{t})q_{0}(z)$ $\alpha_{t}\delta_{z_{1}}(z)+(1-\alpha_{t})q_{0}(z)$ $\alpha_{t}\delta_{z_{1}}(z)+(1-\alpha_{t})q_{0}(z)$
$q_{s|t}(z|Z_{t},z_{1})$ $q_{s|1}(z|z_{1})$ $\frac{q(z_{t}\mid z_{s})\,q(z_{s}\mid z_{1})}{q(z_{t}\mid z_{1})}$ $(1-D_{t})\delta_{z_{t}}(z)+D_{t}\delta_{z_{1}}(z)$
$p_{s|t}(z|z_{t})$ $\alpha_{s}p_{1|t}(z\mid Z_{t})+(1-\alpha_{s})q_{0}(z)$ $\sum_{z_{1}}q_{s|t}(z|Z_{t},z_{1})p_{1|t}(z\mid Z_{t})$ $\delta_{z_{t}}(z)+D_{t}\left(p_{1|t}(z\mid Z_{t})-\delta_{z_{t}}(z)\right)$
Figure 3: Noising and Denoising Processes in Discrete Diffusion and Simple Iterative Denoising
Unlike standard Discrete Diffusion, in Simple Iterative Denoising, an element that is masked (in forward) or unmasked (in backward) can be unmasked (in forward) or remasked (in backward). Note that for the sake of illustration, we use the absence of edge as an absorbing state. In practice, we use three different states for mask SID and mask DDM (mask, edge, non-edge).
Appendix C Generative Graph Modeling: Related Works
A main challenge in generative graph modeling follows from the $n!$ different ways to represent graphs due to node permutation. This has motivated two dominant families of approaches: sequential models, which generate graphs step-by-step, and equivariant models, which preserve equivariance to node permutation by design.
Sequential models
construct graphs auto-regressively, adding nodes, edges, or substructures in a predefined order (You et al., 2018; Shi et al., 2020; Luo et al., 2021; Liao et al., 2019; Kong et al., 2023; Zhao et al., 2024a). To limit the number of possible sequences representing a single graph, many methods adopt a Breadth-First Search (BFS) traversal strategy. In specific domains such as molecular graph generation, canonical formats like SMILES have been used to mitigate permutation issues (Gómez-Bombarelli et al., 2018; Kusner et al., 2017). However, general-purpose canonization strategies (Goyal et al., 2020) fail for large graphs (see experiments in Bergmeister et al. (2024)). Hierarchical models that aggregate subgraphs (Jin et al., 2018, 2020) also fall into this category, though they rely on enumerating a predefined vocabulary of substructures and are thus limited to constrained domains.
Equivariant models
aim to preserve permutation symmetry by producing the same output regardless of node ordering. These models have been instantiated across several generative paradigms, including GANs (Krawczuk et al., 2021; Martinkus et al., 2022), normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020; Liu et al., 2019), and vector-quantized auto-encoders (Boget et al., 2024; Nguyen et al., 2024). More recently, equivariant denoising models in continuous spaces as score-based diffusion (Yang et al., 2019; Jo et al., 2022), diffusion bridges (Jo et al., 2024), and Flow Matching (Eijkelboom et al., 2024) have significantly improved graph generation for small graphs. Despite their success, these continuous methods introduce a mismatch between the continuous noise space and the inherently discrete structure of graphs.
To address this, several works have explored discrete denoising processes. Discrete diffusion models in both discrete time (Haefeli et al., 2022; Vignac et al., 2023) and continuous time (Xu et al., 2024) have demonstrated strong performance in generating small graphs. Discrete flow matching has also been adapted for graphs, with Qin et al. (2024a) showing promising results. However, these discrete models suffer from compounding denoising errors. Our proposed method directly tackles this issue, leading to improved generative performance.
Scalability.